### v2

@@ Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users. Such explanations are essential in enhancing trustworthiness and safety. Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions. This work presents an accountability and explainability architecture implemented for mobile Robot Operating System (ROS)-based robots. The proposed solution consists of two main components. First, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box. The integration of blockchain and generative Artificial Intelligence (AI) solutions ensures the deployment of responsible, transparent, interpretable, and trustworthy agents. The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
@@
Review of blockchain trends for immutable logs, EDRs, XARs, and LLMs.
@@ 1 Introduction
The rising presence of mobile robots alongside humans has increased security and safety concerns. One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios [1]. Therefore, it is essential to ensure that the events performed when developing an action or behavior are reliable and transparent [2]. Equally important is the capacity to render these explanations comprehensible to non-expert users, offering them appropriate, effective, and natural responses to their queries [3, 4].
@@
The most substantial evidence in post-event analysis relates to the logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect any data tampering or unauthorized leakage. Therefore, a trustworthy and secure log system is needed when generating a complete trace of a robot’s behavior [5]. The system must ensure the integrity, confidentiality, and availability of the agent and the data. Additionally, it should possess the capability of incident detection, facilitating investigations, and ensuring compliance, ultimately improving the overall security of the robot.
@@
Regarding log systems vulnerabilities, covering tracks has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. As a result, it is necessary to strengthen system properties against unauthorized modifications [6]. This fact not only includes safeguarding the integrity of log data but also implementing measures to detect and counteract attempts at tampering. Accountability is equally required, enabling the detection and isolation of faulty behaviors and their origin. It must not only provide tamper resistance but also offer verifiability for checking the integrity of log entries against tampering. Moreover, it should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
@@
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, features such as computational power and networking bandwidth can be negatively affected. The inclusion of anti-tampering techniques, such as full disk and communication encryption, may be restrictive due to resource limitations [8]. Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
@@
Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action. For this evidence to be significant, it should belong to a persistent chain of custody, totally reliable and verifiable, main requirements for an auditable and transparent logging system [9].
@@
Blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction, manipulation, or non-permitted editing of data through illicit channels. The nature of this technology helps to avoid tampering risk, safeguarding data [10].
@@
Nevertheless, the potential evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These facts may result in limited usefulness and misalignment with the expectations of non-expert users [11]. Regarding autonomous agents, the raw log messages produced by these systems are usually noisy and possess a semi-structured format. In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [12].
@@
In this particular context, explaining the reasons that make a robot carry out a specific and unexpected action is essential for diagnosing failures in goal achievement. Besides, making these explanations accessible to humans is equally important [13]. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity [14, 15].
Large Language Models (LLMs) have gained significant attention due to their exceptional performance in diverse natural language processing tasks, including translation, coding or text summarization [16, 17, 18]. One approach to improve the results of LLMs for a specific domain is through the use of retrieval-based methods. By retrieving relevant information from a large corpus of text, these models can provide explanations grounded in existing knowledge [19].
The integration of Artificial Intelligence (AI) and blockchain has received significant attention in recent years, with researchers exploring various ways of combining these two areas to provide efficiency, security, reliability and transparency to different domains [20, 21]. Our contribution demonstrates how bringing these technologies together can improve the behavioral analysis of autonomous agents, enhancing their interactions with humans and thus increasing their acceptance.
@@
This work presents an accountability and explainability architecture implemented for ROS-based mobile robots. This solution includes two main components. First, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component. The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events. Both components are connected through an additional component responsible for processing raw data collected by the first one, eliminating non-essential information and building connections and relations between data. These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent. Additionally, we have conducted a thorough evaluation of our solution, including its impact on system performance and the correctness, accuracy, helpfulness and coherence of the generated natural language explanations.
@@ 2 Related Works
In this section, we explore key works related to various domains relevant to our approach. It includes progress in understanding how blockchain ensures the immutability of logs, research on the application of event data recorders in robotics, and significant contributions XAR field. Lastly, we examine the evolution of LLMs to provide natural language explanations.
@@ 2.1 Blockchain for Immutable Logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through hashing and storage of previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process. As a result, data stored in the blockchain is immutable [22, 23].
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24]. While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible and scalable consensus protocol [25, 26].
@@
2.2 Robotic Event Data Recorders
Event Data Recorders (EDRs), often referred to as black boxes, have traditionally played a crucial role in capturing and analyzing events within robotics systems. These recorders support the collection and storage of event data, helping to comprehend system behavior, detecting anomalies, and ensuring safety [35, 36]. Therefore, these solutions are needed in forensic investigations, ethical assessments and transparency regarding autonomous agents [37].
@@ 2.2 Robotic Event Data Recorders
The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38]. These solutions typically include functionalities such as hashing, signing, and encryption [39]. However, to ensure the integrity and authenticity of recorded events, anti-tamper features must also be in place [40].
@@
Using blockchain technology in event data recorders can guarantee the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data, as it becomes difficult for any party to manipulate or tamper with the recorded events without detection [41]. Additionally, blockchain involves decentralized and distributed storage of event data, eliminating the necessity for central authorities or intermediaries and reducing the risk of data loss.
@@
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system proposed in [42], employs permissioned blockchain technology for immutable black box logging in robotic applications. The BBR system combines Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity and immutability. The potential applications of blockchain in robot swarms, including secure communication, data logging, and consensus agreement are discussed in [43]. In [44], a time-segmented consortium blockchain for robotic event registration is proposed, organizing and storing event records efficiently to allow for retrieval and analysis of specific events within specific time frames. Authors in [45] present a tool to bridge Robot Operating System (ROS) and the Ethereum blockchain, evaluating encryption efficiency and stability, transactions response time and packet loss rate. Nevertheless, these works lack capabilities to save integrity proofs in the blockchain at selected intervals, to enhance granularity, or to include service definitions or SCs to improve the verifiability process.
@@ 2.3 Explainable Autonomous Robots
Explainable Artificial Intelligence (XAI) is a research area focused on improving the transparency and interpretability of AI systems, thereby promoting their responsible deployment. The need for explainability in autonomous robots arises from the increasing complexity and autonomy of these systems, making it challenging for users to comprehend their behavior. Explainable Autonomous Robots (XARs) are designed to provide explanations for their actions and decisions to non-expert users. Such explanations enrich aspects like security and the trust level, preventing failures, errors, or conflicts in achieving objectives or undesired interactions [46].
Despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical and effective usability or interpretability in real environments [47]. Therefore, it is a problem that requires special attention, being necessary to explain the performed decisions. This explanation should include alternatives, choices made, expectations of each option, decision criteria, and any unexpected events that may have led to the modification of the initially planned behavior [15, 48].
As described in [49], an explanation regarding XAR is an "abductive reasoning that bridges the gap between the individual explaining, which in this case may be a robot, and the person to whom they are explaining". Self-explanatory capabilities in social robots have been provided through the development of architectures and frameworks [50, 51].
@@
Nevertheless, the essence of explainability regarding XAR lies in effective communication [49]. Progress in natural language processing (NLP) technology through LLMs, based on the estimation of human statements and the relevance and position of the words contained in them, could improve the efficiency of human-robot interactions. This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly benefiting their usefulness and applicability [1].
2.4 Large Language Models
LLMs have recently played a significant role in natural language processing, thanks to replacing recurrent neural networks with self-attention mechanisms [52]. The design and implementation of the first Transformers models constitute the foundation of most LLM architectures. Trained on massive text data, LLMs include strong capabilities in natural language understanding and complex task solving [53].
However, LLMs face two primary challenges. First, they tend to produce a significant amount of hallucinations due to their reliance on internal knowledge (weights) [54]. Second, their large parameter sizes and the associated high update costs, make traditional pre-training and fine-tuning methods unfeasible. As a solution, Retrieval-Augmented Generation (RAG) methods emerge as a solution for LLMs to effectively interact with the external world [55]. This strategy is particularly notable in question-answering systems, which represent the most widely adopted use cases for harnessing the explanatory and reasoning capabilities of LLMs. In this domain, RAG-based approaches are emerging as the standard for developing solutions that seamlessly integrate the advantages of a knowledge base (utilizing a vector store) and generative models.
LLMs display emergent abilities, such as In-Context Learning (ICL), where LLMs make predictions based on contexts augmented with a few examples, helping the model to learn from analogy [56]. This paradigm simplifies the inclusion of human knowledge into LLMs by modifying the prompt templates [57]. ICL does not require altering the model’s parameters or training the model on a specific dataset; instead, it relies on guiding the LLM through the prompting interface to align its results with human values and preferences.
Nevertheless, evaluating LLMs and the impact of previously described approaches poses several challenges due to their complexity and the nature of their outputs. These difficulties can be attributed to several factors, including the lack of suitable evaluation metrics, biases in generated text, and the cost of human assessment [58]. Nevertheless, recently, employing LLMs as evaluators for LLMs-based solutions has been proven to be nearly as effective as human judges [59].
Previous works have demonstrated the potential of using LLMs to generate natural language descriptions of logging systems [60] or robots’ behaviors [61], indicating room for improvements to deliver more accurate and complete interpretations. Nevertheless, to our knowledge, no previous research has investigated the capabilities of LLMs in generating natural language explanations within the context of an accountability approach for record interpretation.
2.5 Artificial Intelligence and Blockchain
XAI and blockchain technology serve as complementary tools to enhance transparency, accountability, trust and acceptance of AI systems. In recent years, there has been a growing recognition of the importance of understanding and interpreting the decisions made by AI systems, particularly within critical domains such as healthcare, finance, and autonomous vehicles [62, 63, 64].
XAI aims to address this need by providing human-understandable explanations for AI predictions and decisions, enabling users to trust and effectively utilize AI systems. However, achieving transparency and interpretability in AI models poses significant challenges, especially when dealing with complex algorithms and large datasets. Through blockchain, a decentralized and immutable ledger can record and verify the decision-making processes of AI models in a transparent and auditable manner.
By integrating blockchain, XAI initiatives can store explanations for AI predictions on a distributed ledger, allowing stakeholders to validate and verify the reasoning behind AI decisions. A method for auditing legal decisions using blockchain technology alongside AI tools is proposed in [21]. Using IPFS and blockchain technology, this approach offers a secure and tamper-proof record to store details, judgements, and reasoning. The use of Generative AI over this data helps to build trust in the system, by analyzing and finding relevant information. Data privacy concerns are also addressed, helping to build a more transparent, accountable, and efficient legal system.
A framework for auditing XAI decisions using blockchain technology is proposed in [28]. This framework includes features for both data storage and verification, with data stored in IPFS and the Ethereum blockchain preserving a unique integrity proof (hash) of the data. This setup allows anyone to verify the integrity of the data, thereby providing transparency and accountability in XAI decision-making processes.
However, these approaches do not directly benefit the autonomous agents’ domain. Instead of relying on IPFS for storing contextual information for explanations, the use of Rosbag2 files, a binary file format, ensures the reproducibility of robots’ behaviors [42]. Rosbags, capable of recording thousands of messages per second, offer a comprehensive and standardized format for capturing robotic data, guaranteeing the recording of each interaction and facilitating accurate analysis. Furthermore, given the inherent indeterminism of LLM algorithms, ensuring the integrity of the messages contained in the Rosbag file through a permissionless blockchain solution offers public transparency and accountability. By storing a chain of integrity proofs of these messages on the blockchain, their immutability is guaranteed, and their verification is simplified in every analysis of the Rosbag file, thereby improving the reliability of the data utilized in the explanation generation process.

### v3

@@ Abstract
The deployment of autonomous agents in environments involving human interaction has increasingly raised security concerns. Consequently, understanding the circumstances behind an event becomes critical, requiring the development of capabilities to justify their behaviors to non-expert users. Such explanations are essential in enhancing trustworthiness and safety. Additionally, they contribute to improving communication, bridging the gap between the agent and the user, thereby improving the effectiveness of their interactions. This work presents an accountability and explainability architecture implemented for mobile Robot Operating System (ROS)-based robots. The proposed solution consists of two main components. First, a black box-like element to provide accountability, featuring anti-tampering properties achieved through blockchain technology. Second, a component in charge of generating natural language explanations by harnessing the capabilities of Large Language Models (LLMs) over the data contained within the previously mentioned black box. The study evaluates the performance of our solution in three different scenarios, each involving autonomous agent navigation functionalities. This evaluation includes a thorough examination of accountability and explainability metrics, demonstrating the effectiveness of our approach in using accountable data from robot actions to obtain coherent, accurate and understandable explanations, even when facing challenges inherent in the use of autonomous agents in real-world scenarios.
@@
Review of blockchain for immutable logs, EDRs, XAI and blockchain, XARs, and LLMs.
@@ 1 Introduction
The rising presence of mobile robots alongside humans has increased concerns about security and trust. One of the main challenges in this context is the ability of robots to justify their actions and decisions, particularly in safety-critical scenarios where their behavior may impact human well-being [1]. Transparency and reliability in these justifications are essential [2], as such properties enable effective human-robot interaction and foster confidence, particularly for non-expert users who require clear and understandable explanations [3, 4]. However, the complexity of robotic decision-making and the need for thorough post-event analysis make achieving these goals particularly difficult.
@@
To address these challenges, accountability must be ensured through a robust, tamper-proof logging system. The most substantial evidence for post-event analysis comes from logging and auditing capabilities, which are essential to mitigate technical issues and provide forensic evidence to detect data tampering or unauthorized breaches. These systems must guarantee the integrity, confidentiality, and availability of data while enabling incident detection, compliance, and forensic investigations [5].
@@
However, covering tracks in log systems has become a common practice for attackers, erasing any traces that could lead to their identification or assisting in the analysis of their actions. As a result, it is necessary to strengthen system properties against unauthorized modifications [6]. This includes safeguarding the integrity of log data and implementing measures to detect and counteract attempts at tampering. Effective accountability solutions must ensure the detection and isolation of faulty behaviors and their origin. Moreover, they should guarantee search accessibility, enabling access to specific records whose integrity can be verified [7].
@@
Integrating accountability solutions into robotic systems often demands a trade-off with system performance. Specifically, computational power and networking bandwidth may be adversely affected [8]. The inclusion of anti-tampering techniques, such as full disk and communication encryption, can be restrictive due to resource limitations [9]. Additionally, high-cost tamper-proof storage devices may prove economically unviable, especially when dealing with continuously generated log data.
@@
Therefore, it is necessary to establish an efficient and verifiable proof of integrity to guarantee the reproducibility and identification of the events that could arise during an autonomous agent action without compromising the robot’s performance. For this evidence to be significant, it should belong to a persistent chain of custody, fully reliable and verifiable, key requirements for an auditable and transparent logging system [10].
@@
To this end, blockchain solutions emerge as a mechanism to ensure data integrity, protection against unauthorized changes, and traceability. The consensus protocols established by all blockchain network members enable the detection and rejection of any destruction or manipulation of data through illicit channels. The nature of this technology mitigates tampering risk, safeguarding data [11].
@@
Nevertheless, while blockchain secures log data, the evidence obtained poses significant challenges for analysis due to its diverse and varied nature, terminology, and the presence of substantial volumes of similar messages unrelated to anomalous events. These factors may result in limited usefulness and misalignment with the expectations of non-expert users [12]. The raw log messages produced by autonomous agents are usually noisy and semi-structured in format. In addition to this, it is necessary to deal with specific features such as large quantities of domain-specific language messages [13]. Addressing this issue requires methods to organize and filter log data, making it accessible and meaningful.
@@
In this context, explaining the reasons behind a robot’s specific and unexpected actions is essential for diagnosing failures and comprehending deviations from intended goals. Equally important is ensuring these explanations are understandable to users [14]. Explainable Artificial Intelligence (XAI) is a research area dedicated to enhancing the transparency and interpretability of Artificial Intelligence (AI) systems to ensure their responsible and ethical deployment. As autonomous robots grow in complexity, understanding and predicting their behavior has become increasingly challenging for non-expert users. This requirement has led to the development of eXplainable Autonomous Robots (XAR), which aims to provide mechanisms for presenting clear and understandable explanations for the actions and decisions of autonomous agents [15]. XAR enhances the transparency and interpretability of these agents’ decision-making processes, often incorporating techniques like natural language explanations to improve clarity and user comprehension [16, 17].
To further improve the effectiveness of XAR systems, Large Language Models (LLMs) excel in diverse natural language processing tasks, including translation, coding, or text summarization [18, 19, 20]. Our approach integrates LLMs with blockchain-secured log data. By using retrieval-based techniques, this solution generates accurate and context-aware natural language explanations based on verifiable information [21]. This integration ensures that explanations are clear, coherent, and grounded in a traceable and reliable context, effectively addressing user comprehension and data integrity concerns.
@@
This work presents an accountability and explainability architecture implemented for Robot Operating System (ROS)-based mobile robots. This solution includes two main components. First, a black box-like component to provide accountability that includes anti-tampering features by distributing integrity proofs from the events recorded by the black box component. The second component generates natural language explanations from data in the black box, including the use of interactive natural language conversations as a means to provide comprehensive explainability for the recorded events. Both components are connected through an additional module responsible for processing raw data collected by the first one, eliminating non-essential information, and building connections and relations between data. These functionalities release software developers from the need to make calls to the logging Application Programming Interface (API) to build explanations for the code generated and the behaviors developed by the autonomous agent. To validate our approach, we conducted a comprehensive evaluation, assessing the architecture’s impact on system performance and analyzing the quality of the generated explanations. The evaluation considered key metrics, including correctness, accuracy, helpfulness, coherence, and user satisfaction. Our results demonstrate that the architecture effectively balances accountability and explainability without compromising the robot’s performance, providing a scalable and reliable framework for real-world applications.
@@ 2 Related Works
In this section, we explore key works related to various domains relevant to our approach. This includes advances in how blockchain ensures the immutability of logs, research on applying Event Data Recorders (EDRs) in robotics, significant contributions to integrating XAI with blockchain, and XAR. Finally, we examine the evolution of LLMs for generating natural language explanations and their advancements within robotics.
@@ 2.1 Blockchain for Immutable Logs
Blockchain is a decentralized solution that eliminates the necessity for trust in a central authority through its immutable and distributed ledger, composed of timestamped transaction blocks. These blocks are linked through hashing and the storage of previous block content, ensuring the integrity of transactions. Any attempt to alter a transaction in one block requires the modification of all subsequent blocks, resulting in a computationally expensive process. As a result, data stored in the blockchain is immutable [22, 23].
There are two primary types of blockchains based on access controls: permissionless blockchains allow open access and public transaction visibility, while permissioned blockchains limit participation and transaction visibility. Nevertheless, closed networks may pose security risks, as authorized users could recalculate subsequent blocks in the event of a block modification [24]. While permissionless blockchains have traditionally related to scalability and throughput challenges, approaches like Ethereum have transitioned to Proof of Stake (PoS), a more energy-efficient, secure, accessible, and scalable consensus protocol [25, 26].
@@
Our architecture uses the immutability of blockchain to store evidence of the integrity of recorded data, thus providing real-time guarantees against manipulation and ensuring that the explanations generated by our system are always based on non-manipulated data.
2.2 Blockchain for Event Data Recorders in Robotics
@@
EDRs, or black boxes, are essential for capturing and analyzing events in robotic systems. These recorders help to comprehend system behavior, detect anomalies, and ensure safety [35, 36]. Therefore, these solutions are needed in forensic investigations and ethical assessments regarding autonomous agents [37]. The integration of black box modules to enhance robots’ introspection capabilities contributes to improving system transparency by attributing responsibilities [38, 9]. These solutions typically include functionalities such as hashing, signing, and encryption [39], while ensuring event integrity through anti-tamper mechanisms [40].
@@
Using blockchain technology in EDRs ensures the integrity of a transparent and auditable record of events. These features enhance accountability and trust in the recorded data by preventing tampering [41]. Decentralized and distributed storage further eliminates reliance on central authorities and mitigates data loss risks.
@@
Further research explores the application of blockchain technology to improve robotic capabilities, particularly event data recording and management. The Black Block Recorder (BBR) system uses permissioned blockchain for immutable logging, combining Rosbag2 and Distributed Ledger Technologies (DLTs) to ensure data integrity [42]. Blockchain applications in robot swarms, including secure communication, data logging, and consensus agreements, are discussed in [43]. In [44], a time-segmented consortium blockchain for robotic event registration is proposed, enabling efficient event organization and retrieval within specific time frames. Authors in [45] present a tool to bridge ROS and the Ethereum blockchain, evaluating encryption efficiency and stability, transaction response time, and packet loss rate. Nevertheless, these works lack the capability to save integrity proofs in the blockchain at selected intervals, enhance granularity, or include service definitions or SCs to improve the verifiability process.
2.3 Explainable Artificial Intelligence and Blockchain
XAI aims to align AI decisions with human cognitive expectations by employing techniques such as contextual relationships, contrastive explanations, and methods like counterfactual reasoning and interpretable models that enhance interpretability and transparency [46, 47]. Insights from social sciences show people tend to interpret explanations more positively when they reflect human-like reasoning, emphasizing intentions, goals, and contextual factors rather than purely technical or statistical descriptions [48]. By combining XAI with blockchain transparent record-keeping, explanations can be retraced and validated, ensuring both interpretability and accountability in AI-driven systems [49, 50].
The benefits of this integration are remarkable in various domains. In healthcare, XAI and blockchain have been combined for secure, transparent patient data management [51], logical disease prediction, and immutable medical diagnostics records [52]. Similarly, in the legal domain applications, blockchain and XAI support tamper-proof decision auditing, using IPFS and Ethereum for cryptographic proofs that enable data verification and accountability [53, 54, 28]. Privacy-preserving systems with homomorphic encryption and semantic ontologies further ensure data protection while maintaining explainability [55].
Regarding autonomous agents like unmanned aerial vehicles or autonomous vehicles, blockchain and XAI ensure secure and transparent decision-making [56]. This integration enhances collaboration, even in complex and time-critical scenarios [57]. Decentralized edge models for collaborative AI, leveraging blockchain for secure communication and real-time data sharing, further optimize resources and enhance reliability in these environments [58].
Our work introduces Rosbag2 files as an alternative to IPFS or distributed file systems for storing contextual data in autonomous robotic systems. This approach supports reproducibility and enables a more precise analysis of robotic behaviors [42]. Rosbag files provide a standardized format for recording large volumes of robotic data, facilitating comprehensive interaction capture for improved analysis. To address the indeterminism of LLM algorithms, we propose using a permissionless blockchain to store integrity proofs of Rosbag messages, ensuring immutability, simplifying verification, and enhancing data reliability for explanation generation.
@@ 2.4 Explainable Autonomous Robots
XARs are designed to provide explanations for their actions and decisions to non-expert users. Such explanations enhance security and trust, helping to prevent failures, errors, conflicts, and undesired interactions while achieving objectives [59, 60]. However, despite the growing demand for explainability in autonomous agents, their use reveals a lack of practical usability and effective interpretability in real environments [61].
Effective explanations in XARs should detail alternatives, the choices made, expectations for each option, decision criteria, and any unexpected events that may have led to changes in the planned behavior [17, 62]. These explanations often rely on abductive reasoning, which bridges the gap between the robot’s internal decision-making process and the user’s understanding [63]. To support this, frameworks have been developed that enable robots to articulate their internal states and decisions, enhancing collaboration and improving communication with humans [64].
Adaptive algorithms and needs-based motivational systems allow robots to provide real-time, context-aware explanations tailored to diverse audiences [65, 66]. Behavior Trees (BTs) hierarchically decompose tasks into goals and subgoals, making complex behaviors easier to understand and adaptable to real-time changes [67]. However, achieving a balance between simplicity and adequacy is essential when making complex information easy to understand. By focusing on the most relevant details and filtering non-essential information, explanations foster trust and collaboration by aligning with human cognitive capacities [68]. For this purpose, our architecture incorporates a log curation module that processes raw robotic data, filtering it to retain only meaningful, high-relevance information, thus delivering clearer and more concise explanations directly tied to events.
@@
In addition to context-aware explanations, the core of explainability in XAR lies in effective communication [63]. Progress in Natural Language Processing (NLP) and LLMs can improve the efficiency of human-robot interactions. This enhancement can create a sense that these systems truly comprehend and respond to individuals meaningfully, significantly improving their usefulness and applicability [1].
2.5 Large Language Models in Robotics
LLMs, enabled by self-attention mechanisms, have advanced NLP, improving task-solving capabilities and enabling more effective human-robot interaction [69, 70, 71]. However, LLMs face challenges such as hallucinations, which arise due to reliance on internal knowledge, and the high update costs associated with large parameter sizes, making traditional pre-training and fine-tuning methods impractical. Retrieval-Augmented Generation (RAG) methods address these issues by enabling LLMs to interact effectively with external sources [72, 73]. This approach is particularly advantageous in robotics, where real-time and accurate contextual understanding is essential for complex decision-making and task execution.
Explainability and transparency have become critical aspects of deploying LLMs in autonomous agents. These models enable robots to translate high-level commands into actionable plans, improving performance in complex environments [74, 75, 76]. When integrated with reward decomposition frameworks, LLMs can generate intuitive explanations, linking actions to object-specific properties, thereby reducing ambiguity [77]. LLMs can also generate natural language descriptions of robot behaviors and logs, highlighting opportunities for improvement in accuracy and completeness [78, 79].
Existing solutions, however, often rely on static models or predefined data, limiting their ability to handle dynamic environments. Our approach addresses these limitations by integrating blockchain-supported accountability with curated logging of relevant robotic data. This framework provides robust context for LLM-generated explanations, minimizing hallucinations and ensuring alignment with the robot’s actions and interactions within its environment.
Evaluating LLMs in robotics presents additional challenges, including the absence of suitable metrics, biases in generated text, and the high cost of human evaluations [80]. While human evaluations remain the gold standard for capturing subjective insights [48], they are resource-intensive and lack scalability [81]. LLM-based evaluations offer a scalable alternative, achieving over 80% alignment with human judgments while delivering explainable outputs [82]. A hybrid framework combining human assessments for nuanced insights and LLM evaluations for efficiency and repeatability provides a comprehensive and scalable solution for evaluating LLM outputs.